# On the attribution of weather events to climate change using a fit to extreme value distributions

### v2

@@ 3.2 Siberian Heatwave
We now explore other facets of the attribution of the Siberian Heatwave in the summer of 2020 to anthropogenic change, notwithstanding the foregoing finding that the interpretation of $\alpha$ is complicated by the presence of natural variability. We focus specifically on the CC2021 analysis of the June daily maximum temperatures because it allows us to address the question of whether the data justify the use of non-stationary (GMST-dependent) extreme value distributions.
@@
While it is understood based on theoretical considerations that a GEV is the appropriate representation of block-maxima of data Coles et al. (2001), it is not obvious for a given problem what minimum block size justifies using a GEV, and specifically if a year of daily maximum data suffices. This is especially an issue given that the temperature is auto-correlated and, therefore, more data are needed per averaging block to justify the use of a GEV. It is, therefore, worthwhile to check how well a more standard (say normal) distribution performs in fitting the data.
@@
A cumulative distribution estimate of the highest June daily-maximum temperature observed each year between 1926–2020 at the Verkhoyansk station is shown in Fig. 2a along with a 95% confidence interval of the CDF obtained by bootstrapping. In the first step of our analysis, we consider cumulative distribution function fits to the observed distribution, assuming the data follows either a GEV or normal distribution. The fitted distributions are specified here to be stationary (by setting $\alpha=0$ and $\mu=\mu_{0}$ in Eqns. 14). Fits are obtained by maximizing the likelihood of the data, and the results are shown in Fig. 2a. Both distributions generally fit the observations well, with the normal fit completely within the 95% range of the observed distribution, and with the GEV fit being good except at the highest temperature, where it slightly underestimates the probability of obtaining the highest temperatures. There is only one data point above 35 °C, so the empirical CDF is poorly constrained there in any case. Interestingly, the normal distribution and GEV CDFs look similar and lie within the confidence intervals defined for the empirical CDF, despite the normal distribution having fewer parameters than the GEV. It is worth noting that a GEV and a normal distribution may assign very different probabilities to extreme events even when the two distributions provide an adequate fit to the observed CDF.
@@
The quality of fits seems to suggest that an extreme value distribution may not be needed as a normal distribution does just as well. Similarly, a non-stationary distribution with $\alpha\neq 0$ may not be necessary given that the stationary distributions seem to provide a satisfactory fit to the empirical cumulative distributions, and we further examine this next. While CDFs are commonly used to differentiate between distribution functions, as in the Kolmogorov-Smirnov test, plots of the PDFs (Supplementary Fig. SI-1) for the three distributions (normal, GEV and empirical) also suggest that the normal distribution aligns well with the observed data (and mostly that the data may not be sufficient to well-constrain the distribution). An application of Occam’s razor principle would, therefore, suggest that the simpler Gaussian distribution should be used rather than a GEV if there is not a significant improvement in the distribution fit with the GEV. This is also discussed by Philip et al. (2020) who mention explicitly that the extreme value distributions may not outperform a simpler (Gaussian) distribution in some cases. Otto et al. (2018) concluded that a Gaussian distribution is justified because the events in question are not very rare. CC2021 found that a Gaussian distribution fits the data well but proceeded to use a GEV for the bulk of the analysis. These results all question the need for extreme value distribution functions when it comes to attribution analysis.
@@
Figure 2: Siberian Heatwave: (a) Testing the need for a GEV fit. The red curve indicates the empirical CDF of the highest daily maximum temperature for each year for the Verkhoyansk station data. The purple and blue lines denote fits by a GEV (with a constant location parameter) and by a normal distribution, respectively. The light and dark green shadings correspond to 90th and 95th percentile confidence intervals (section 22.3). (b) Justification for using a GMST-dependent mean. Box plots of the deviance statistic between a GEV distribution without a GMST-dependent mean relative to a GEV distribution with a GMST-dependent mean using the Siberian heatwave data with 5,000 bootstrap resamples. The edges of the boxes correspond to the first and third quartiles. The deviance statistic for the station observations is shown in green as the leftmost box. The deviance statistic was also computed for each of the 35 members from the CESM Large Ensemble (blue). The 95th percentile significance level from a $\chi_{k=1}^{2}$ distribution with one degree of freedom is shown by the horizontal red line, corresponding to the addition of one degree of freedom by including $\alpha$ in the fitted model.
@@
We now proceed to examine the role of GMST in the fit to a GEV distribution (eq. 4) in this particular application, following CC2021. The value of $\alpha$ and its uncertainty range for the Siberian Heatwave do not seem to be explicitly discussed by CC2021. A linear regression of GMST anomalies against daily max temperatures gives a slope of 1.1 C per C (Fig. 3a), where the positive value indicates that the annual maximum daily temperatures at Verkhoyansk scale with GMST. The squared Pearson correlation, however, is only $r^{2}=0.01$, with a p-value of 0.2, indicating that GMST is only a weakly significant indicator of trends in Siberian temperature maxima.
@@
Fig. 3b shows the maximum likelihood fit of a nonstationary GEV, including $\alpha$, for multiple bootstrapping samples as specified in the Methods section. The estimate for the observations $\alpha$ without resampling in the observations is 1.8 C per C, similar to the slope of 1.1 obtained from the linear regression. The [2.5%, 97.5%] confidence interval is [$-1.2$, 4.8], while the 5% value is $-0.7$ (Fig. 3b; red plots). Both indicate that $\alpha$ is not significantly different from zero. While CC2021 did not provide an uncertainty range for $\alpha$, they mention that the change in intensity of Siberian heatwaves (at a fixed probability) is estimated from their estimated PDF as 1.04 C (with a 95% confidence interval from 0.35 to 3.4). There is no straightforward correspondence between this uncertainty range and that we calculate for $\alpha$, unfortunately. We do note that the constraint that $|\xi|<0.4$ imposed by CC2021 likely reduces their uncertainty range relative to the analysis we present here. The solution for all GEV parameters for the observations is given in Supplementary Fig. SI-2.
@@
The deviance statistics, $D$, (section 22.4) indicate that for the observations (Fig. 2b, leftmost bar), the addition of one more model parameter, $\alpha$, representing the GMST dependence, is justified as the value is above the threshold indicated by the red horizontal line. Yet the deviance statistic uncertainty range includes a large range that is below the significance line, which seems appropriate given that the uncertainty range for $\alpha$ includes negative values.
@@
To provide context for the interpretation of the observational results, we also examine output from the CESM Large Ensemble, which has time-varying anthropogenic forcing. We find that all ensemble members within the CESM Large Ensemble indicate a positive value of $\alpha$, with a mean value of 2.60 C per C (Fig. 3b; blue plots). Similarly, 33 of 35 ensemble members give a deviance statistic that supports the use of a non-stationary component (that is, a GMST-dependence) in the GEV statistic (Fig. 2b).
@@
The foremost reason that the Large Ensemble suggests a stronger link between Siberian extremes and GMST (that is, a larger $\alpha$) may be that the LE, on average, simulates an Arctic-wide summer warming rate of 1.8 C per C warming in GMST, whereas an observational analysis only indicates a rate of 1.3 C Arctic JJA warming per C in GMST (using GISTEMP for both; Fig. 5a,b). One expects heatwaves to be strongly influenced by the JJA climatology, and Fig. 5c shows that different ensemble members, corresponding to different realizations of natural variability, show very different dependencies of the Arctic JJA climatology on GMST. It is difficult, therefore, to estimate from the single realization corresponding to the observed record to what degree the Arctic JJA climatology is affected by anthropogenic climate change versus natural variability, adding to the uncertainty in attributing Siberian heatwaves. To address that, the WWA protocol includes a step where the same analysis is applied to climate models, and results are reported from both sources.
@@
Unsurprisingly, if data from multiple ensemble members are analyzed simultaneously as a single dataset, which mimics the availability of more data, the distribution of the GEV parameters narrows, generally converging around the mean value of the parameters across individual ensemble member fits (Fig. 4). The 95% confidence range for $\alpha$ (red asterisks in panel a) decreases from a large range of [1,4] (C per C) with one ensemble member to a smaller [2,3] range that may allow for a more reliable attribution with 10 ensemble members. Similar behavior is seen for the other GEV parameters. This implies that with a sufficiently long record, one can reduce the uncertainty in the GEV parameters, as expected, but that might require much more data than is expected to be available in the historical record anytime soon.
@@
The overall result for the GMST dependence for the Siberian Heatwave analysis, which is the base of the attribution to anthropogenic climate change, seems superficially consistent with the results found from CC2021. However, the important caveats are that (1) we found in the previous subsection that internal variability can lead to GMST dependence that biases the value of $\alpha$ and hence the attribution results (Fig. 1a); (2) a GEV fit may, a priori, make sense to fit to annual daily temperature extremes being a block maxima time series. However, a normal distribution with $\alpha=0$ does not appear to be ruled out for the observations (Fig. 2a). This may be a result of both the relatively short observational record and the autocorrelated nature of the temperature extremes which results in one year not being a sufficiently long block to justify a GEV fit. Still, a non-zero $\alpha$ is found to meaningfully improve the statistical fit to the CESM Large Ensemble Siberian extremes based on deviance statistics; (3) the uncertainty range for $\alpha$, not explicitly mentioned in the original WWA analysis, includes negative values (Fig. 3b).
@@
Figure 5: Natural variability and the attribution of Siberian heatwaves. (a) A scatter plot of the annual average JJA Arctic (north of $70^{\circ}$ N) surface temperatures against the 4-year running mean of GISTEMP annual GMST. (b) Same, for the 35 members of the CESM Large Ensemble over the period 1926–2019. (c) PDF of the slopes of annual average JJA Arctic surface temperatures against the 4-year running mean of annual GMST within each member of the CESM Large Ensemble with 5,000 bootstrap resamples (red). The same distribution is indicated for GISTEMP observations in blue. The vertical red line indicates the average slope across the CESM Large Ensemble, and the vertical blue line indicates the slope determined from observations.
3.3 Australian Bushfire
@@
We now consider the analysis of possible connections between the hot weather that may have contributed to the Australian Bushfires of the summer of 2020–2021 and anthropogenic climate change, using both observations and the CESM Large Ensemble, following the observational analysis of OK2021. We again find that a normal distribution and a GEV distribution (both with a constant mean/location parameter) fit the 7-day moving mean of daily temperature maxima in the observations without a GMST dependence well (Fig. 6a). In this case, a normal distribution fit to the empirical distribution has a mean absolute error (MAE) of 0.01, equal to the MAE for the GEV fit. Plots of the PDFs (Supplementary Fig. SI-3) for the three distributions (normal, GEV, and empirical) again suggest that the normal distribution aligns well with the observed data (to some degree due to the small number of data which does not constrain the PDF very well). Given that the normal distribution has one fewer parameter, it seems to offer a simpler approach to representing the data. The deviance statistic for the observations does suggest a significant GMST component (leftmost bar in Fig. 6b). However, given the possibility that this component is driven by natural variability rather than by anthropogenic change (section 33.1, Fig. 1a), one must conclude again that even if $\alpha$ is justified statistically, it might represent the effects of natural variability rather than anthropogenic climate change. In that case, it would be best not to use it for the purpose of attribution analysis, of course. The deviance statistic test for the CESM Large Ensemble shows that only five of the 35 ensemble members justify adding GMST dependence to the distribution parameters (Fig. 6b). This raises the question of whether the influence of GMST, and therefore of anthropogenic climate change in the WWA formulations, is a necessary part of the statistical description of extreme temperatures in Australia as represented in the CESM Large Ensemble.
@@
Figure 6: Australian Bushfire: (a) Testing the justification for a GEV. The red curve indicates the empirical CDF of the highest annual 7-day running mean of daily maximum temperature over the area defined in van Oldenborgh et al. (2021) for observations. The purple and blue lines denote GEV (with a constant mean), and normal distribution fits, respectively. The light and dark green shadings correspond to 90th and 95th percentile confidence intervals. (b) Justification for using a GMST-dependent mean. Box plots of the deviance statistic between a GEV without a GMST component and a GEV with a GMST component using the Australia Bushfire example with 5,000 bootstrap resamples. The edges of the boxes correspond to the first and third quartiles. The deviance statistic was computed for each of the 35 members from the CESM Large Ensemble (blue) and compared with the 95th percentile significance level from a $\chi^{2}$ distribution with one degree of freedom (red line). A box plot for the deviance statistic for observations is shown in green.
@@
The regression analysis of extreme temperatures and GMST, (Fig. 7a) as well as the range of $\alpha$ values estimated from observations (Fig. 7b; red) both show a GMST dependence near 2 C change in extreme temperatures per degree increase in GMST. As mentioned above, a sensible null hypothesis for heatwaves in a warming climate is that the extreme values shift with the mean warming (Tziperman, 2022). The GMST dependence calculated here suggests that Australian heatwaves increase at nearly twice the rate of the GMST. Such a strong response requires a physical mechanism. Some candidates are (1) the fact that land heats up more quickly than the ocean, so GMST always tends to lag behind land surface temperatures; (2) soil-drying can lead to further increases in warming over land as less water is available for evaporative cooling, etc. It is not clear that these can lead to a doubling of the GMST effect and further exploration of this is beyond the scope here. The alternative is that this is a result of natural variability (likely ENSO) affecting both Australian heat extremes and the GMST, consistent with Fig. 1a. Again, this suggests that using this value of $\alpha$ for attribution may lead to unreliable results. Finally, Fig. 7b (blue) shows a large range of GMST-dependencies for different ensemble members in the CESM Large Ensemble, again highlighting that a large component of this dependence most likely represents natural variability rather than anthropogenic change. The solution for all GEV parameters for the observations is given in Supplementary Fig. SI-4. How the GEV parameters scale with the number of CESM Large Ensemble members used is indicated in Supplementary Fig. SI-5.
@@
Figure 7: Australian Bushfire: distribution of $\alpha$ values. (a) Scatter plot of the highest annual 7-day running mean of daily maximum temperatures against GMST (following van Oldenborgh et al., 2021). The red line indicates a linear regression fit to the data, and the blue line indicates a linear fit using the location parameter and $\alpha$ derived from the GEV maximum likelihood estimate. (b; red) PDF of the GEV-fit $\alpha$ distributions from observations, using bootstrap resampling. The $\alpha$ value from the observations without resampling is shown by the red line (at 1.7) (b; blue) Effect of the natural variability and number of data from a model large ensemble on the uncertainty in $\alpha$ for the Australian Bushfire example. Shown is a histogram of the GEV-fit $\alpha$ PDFs derived from the $\alpha$ value from each ensemble member with no resampling. The average $\alpha$ value across the Large Ensemble is indicated with the blue line.

### v3

@@ 3.2 Siberian Heatwave
We now explore other facets of the attribution of the Siberian Heatwave in the summer of 2020 to anthropogenic change, notwithstanding the foregoing finding that the interpretation of $\alpha$ is complicated by the presence of internal variability. We focus specifically on the CC2021 analysis of the June daily maximum temperatures because it allows us to address the question of whether the data justify the use of non-stationary (GMST-dependent) extreme value distributions.
@@
While it is understood based on theoretical considerations that a GEV is the appropriate representation of block-maxima of data Coles et al. (2001), it is not obvious for a given problem what minimum block size justifies using a GEV, and specifically if a year of daily maximum data suffices. This is especially an issue given that the temperature is auto-correlated and, therefore, more data are needed per averaging block to justify the use of a GEV. It is, therefore, worthwhile to check how well other standard distributions perform in fitting the data.
@@
A cumulative distribution estimate of the highest June daily-maximum temperature observed each year between 1926–2020 at the Verkhoyansk station is shown in Fig. 2a along with a 95% confidence interval of the CDF obtained by bootstrapping. In the first step of our analysis, we consider cumulative distribution function fits to the observed distribution, assuming the data follows either a GEV or normal distribution. The fitted distributions are specified here to be stationary (by setting $\alpha=0$ and $\mu=\mu_{0}$ in Eqns. 14). Fits are obtained by maximizing the likelihood of the data, and the results are shown in Fig. 2a. Both distributions generally fit the observations well, with the normal fit completely within the 95% range of the observed distribution, and with the GEV fit being good except at the highest temperature, where it slightly underestimates the probability of obtaining the highest temperatures. There is only one data point above 35 °C, so the empirical CDF is poorly constrained there in any case. Interestingly, the normal distribution and GEV CDFs look similar and lie within the confidence intervals defined for the empirical CDF, despite the normal distribution having fewer parameters than the GEV. It is worth noting that a GEV and a normal distribution may assign very different probabilities to extreme events, even when the two distributions provide an adequate fit to the observed CDF. These modeling decisions are therefore of critical importance for attribution studies, which usually focus as much or more on estimation of the exceedance probability and return period of an event — and, consequently, on the changing likelihood and fraction of attributable risk of an event — as they do on the change in intensity. It is therefore useful to consider alternate distributional forms and, in many ways, the normal is the simplest and often a useful default.
@@
The quality of fits indicates that an extreme value distribution is not empirically a better choice than a normal distribution for fitting the data. Similarly, a non-stationary distribution with $\alpha\neq 0$ may not be necessary given that the stationary distributions seem to provide a satisfactory fit to the empirical cumulative distributions, and we further examine this next. CDFs are commonly used to differentiate between distribution functions, and the results of a Kolmogorov-Smirnov test (Supplementary Fig. SI-1 for normal, GEV and empirical distributions) also suggests that the normal distribution is consistent with the observed data. An application of Occam’s razor principle suggests that the simpler normal distribution should be used rather than a GEV if there is not a significant improvement in the distributional fit with the GEV. This is also discussed by Philip et al. (2020) who mention explicitly that the extreme value distributions may not outperform a simpler (normal) distribution in some cases. Otto et al. (2018) opted to use a normal distribution in lieu of a GPD for attribution analysis of a cold spell in Peru because the normal “uses all data, rather than the GPD, which samples only the cold tail”. CC2021 found that a normal distribution fits the data well but proceeded to use a GEV for the bulk of the analysis. These results question the need for extreme value distribution functions when it comes to attribution analysis. One might want to use proper selection criteria (the Akaike Information Criterion, AIC, or deviance statistic or similar) to rigorously decide which of the models performs best.
@@
Figure 2: Siberian Heatwave: (a) Examining the fit of a GEV and normal distribution to Verkhoyansk station data. The red curve indicates the empirical CDF of the highest daily maximum temperature for each year for the Verkhoyansk station data. The purple and blue lines denote fits by a GEV (with a constant location parameter) and by a normal distribution, respectively. The light and dark green shadings correspond to 90th and 95th percentile confidence intervals (section 22.3). (b) Justification for using a GMST-dependent mean. Box plots of the deviance statistic between a GEV distribution without a GMST-dependent mean relative to a GEV distribution with a GMST-dependent mean using the Siberian heatwave data with 5,000 bootstrap resamples. The edges of the boxes correspond to the first and third quartiles. The deviance statistic for the station observations is shown in green as the leftmost box. The deviance statistic was also computed for each of the 35 members from the CESM Large Ensemble (blue). The 95th percentile significance level from a $\chi_{k=1}^{2}$ distribution with one degree of freedom is shown by the horizontal red line, corresponding to the addition of one degree of freedom by including $\alpha$ in the fitted model.
@@
We now proceed to examine the role of GMST in the fit to a GEV distribution (eq. 4) in this particular application, following CC2021. A linear regression of GMST anomalies against daily max temperatures gives a slope of 1.1 C per C (Fig. 3a), where this value indicates that the annual maximum daily temperatures at Verkhoyansk increase slightly more quickly than GMST. The squared Pearson correlation, however, is only $r^{2}=0.01$, with a p-value of 0.2, indicating that GMST is only a weakly significant indicator of trends in Siberian temperature maxima.
@@
Fig. 3b shows the maximum likelihood fit of a nonstationary GEV, including $\alpha$, for multiple bootstrapping samples as specified in the Methods section. The values of $\alpha$ estimated from the observations is 1.8 C per C, larger than the slope of 1.1 obtained from the linear regression. This discrepancy between $\alpha$ and the slope suggests that the conclusions from an attribution analysis can be highly sensitive to the choice of the fitted distribution, again underscoring the importance of model selection. The [2.5%, 97.5%] confidence interval is [$-1.2$, 4.8], while the 5% value is $-0.7$ (Fig. 3b; red plots). Both indicate that $\alpha$ is not significantly different from zero. It is noteworthy that the red distribution in Fig. 3b is bimodal, presumably because some of the bootstrapped samples included the 2020 event and some did not. This suggests that the fitted distribution is very sensitive to the inclusion or exclusion of the event (again consistent with Miralles and Davison (2023)); that is, $\alpha$ would be significantly non-zero if the event is included, but is not significant when it is omitted. While CC2021 did not provide an uncertainty range for $\alpha$, they note that the change in intensity of Siberian heatwaves at a fixed probability is estimated from their inferred PDF as 1.04 C (with a 95% confidence interval from 0.35 to 3.4). $\alpha$ is directly related to the change in intensity: the change in intensity is $k*\alpha$, where $k$ is the observed GMST change between the two periods of interest, usually 1.2C. Assuming a GMST change of 1.2C, we infer the value of $\alpha$ in CC2021 is 1.3 with a 95% confidence interval from 0.42 to 4.1. The central estimate of $\alpha$ is similar to our results but entails a smaller uncertainty range, likely due to our decision to include the 38C event in our bootstrap resampling, but which was not done in CC2021. The bootstrapped distribution for all GEV parameters for the observations is given in Supplementary Fig. SI-2.
@@
The deviance statistics, $D$, (section 22.4) indicate that the addition of a model parameter, $\alpha$, representing the GMST dependence, is justified in the case of the Verkhoyansk observations, ((Fig. 2b, leftmost bar is above the threshold indicated by the red horizontal line). Yet the deviance statistic uncertainty range includes a large range that is below the significance line, which seems appropriate given that the uncertainty range for $\alpha$ includes negative values.
@@
To provide context for the interpretation of the Verkhoyansk observational results, we also examine output from the CESM Large Ensemble, which includes time-varying anthropogenic forcing, at the grid point containing the Verkhoyansk station. We find that all ensemble members within the CESM Large Ensemble indicate a positive value of $\alpha$, with a mean value of 2.60 C per C (Fig. 3b; blue plots). Similarly, 33 of 35 ensemble members give a deviance statistic that supports the use of a non-stationary component (that is, a GMST-dependence) in the GEV statistic (Fig. 2b).
@@
The fact that the LE results indicate a stronger link between Siberian extremes and GMST (that is, a larger $\alpha$) is consistent with the LE, on average, simulating an Arctic-wide summer warming rate of 1.8 C per C warming in GMST, whereas an observational analysis only indicates a rate of 1.3 C Arctic JJA warming per C in GMST (using GISTEMP for both; Fig. 4a,b). This greater amplification of Arctic warming may be related to the fact that CESM1 has a relatively high climate sensitivity of 4.1 K Gettelman et al. (2019). The WWA approach leverages a range of climate models so that results are not dependent entirely on one GCM. That said, one expects heatwaves to be strongly influenced by the JJA climatology, and Fig. 4c shows that different ensemble members, corresponding to different realizations of internal variability, show very different dependencies of the Arctic JJA climatology on GMST. It is difficult, therefore, to estimate from the single realization corresponding to the observed record to what degree the Arctic JJA climatology is affected by ACC versus internal variability, adding to the uncertainty in attributing Siberian heatwaves. To address that, the WWA protocol includes a step where the same analysis is applied to climate models, and results are reported from both sources.
@@
Unsurprisingly, if data from multiple ensemble members are analyzed simultaneously as a single dataset, which mimics the availability of more data, the distribution of the GEV parameters narrows, generally converging around the mean value of the parameters across individual ensemble member fits (Fig. 5). The 95% confidence range for $\alpha$ (red asterisks in panel a) decreases from a large range of [1,4] (C per C) with one ensemble member to a smaller [2,3] range with 10 ensemble members. Similar behavior is seen for the other GEV parameters. This implies that with a sufficiently long record, one can reduce the uncertainty in the GEV parameters, as expected, but that might require much more data than is expected to be available in the historical record anytime soon.
@@
The overall result for the GMST dependence for the Siberian Heatwave analysis, which is the basis of the attribution to ACC, is largely consistent with the results found from CC2021 but with three important caveats. First, we found in the previous subsection that internal variability can lead to GMST dependence that biases the value of $\alpha$ and hence the attribution results (Fig. 1a). Second, whereas a GEV fit a priori make sense to fit to block maxima of daily temperature extremes, a normal distribution with $\alpha=0$ does not appear to be ruled out by the observations (Fig. 2a). This may be a result of both the relatively short observational record and the autocorrelated nature of the temperature extremes. Still, a non-zero $\alpha$ is found to meaningfully improve the statistical fit to the CESM Large Ensemble Siberian extremes based on deviance statistics. Finally, the uncertainty range for $\alpha$, which was not explicitly evaluated in the original WWA analysis, includes negative values (Fig. 3b).
Figure 4: Internal variability and the attribution of Siberian heatwaves. (a) A scatter plot of the annual average JJA Arctic (north of $70^{\circ}$ N) surface temperatures against the 4-year running mean of GISTEMP annual GMST. (b) Same, for the 35 members of the CESM Large Ensemble over the period 1926–2019. (c) PDF of the slopes of annual average JJA Arctic surface temperatures against the 4-year running mean of annual GMST within each member of the CESM Large Ensemble with 5,000 bootstrap resamples (red). The same distribution is indicated for GISTEMP observations in blue. The vertical red line indicates the average slope across the CESM Large Ensemble, and the vertical blue line indicates the slope determined from observations.
@@
3.3 Australian Bushfire Heatwave
@@
We now consider the analysis of possible connections between the hot weather that may have contributed to the Australian Bushfires of the summer of 2020–2021 and ACC, using both observations and the CESM Large Ensemble, following the observational analysis of OK2021. We again find that a normal distribution and a GEV distribution (both with a constant mean/location parameter) fit the 7-day moving mean of daily temperature maxima in the observations without a GMST dependence well (Fig. 6a). In this case, a normal distribution fit to the empirical distribution has a mean absolute error (MAE) of 0.01, equal to the MAE for the GEV fit. Supplementary Fig. SI-3 shows results for normal, GEV, and empirical distributions that again suggest that the normal distribution aligns well with the observed data. The alignment may reflect that the small number of available data do not constrain the PDF very well. Given that the normal distribution has one less parameter, it seems to offer a simpler approach to representing the data. The deviance statistic for the observations does suggest a significant GMST component (leftmost bar in Fig. 6b). However, given the possibility that this component is driven by internal variability rather than by anthropogenic change (section 33.1, Fig. 1a), one must conclude again that even if $\alpha$ significantly differs from zero, this might reflect a bias introduced by the effects of internal variability. This underscores the need to include covariates representing relevant internal variability based on our physical understanding. The deviance statistic test for the CESM Large Ensemble shows that in only5 of the 35 ensemble members is the extra complexity of adding GMST dependence to the distribution justified (Fig. 6b). This raises the question of whether the influence of GMST, and therefore of ACC in the WWA formulation, is a necessary part of the statistical description of extreme temperatures in Australia as represented in the CESM Large Ensemble. It should be noted that our results relate only to one climate model that, of course, does not necessarily represent all the relevant physical processes at work in the real world. The use of multiple climate models in the WWA analysis helps ensure a wider range of process representations are included.
@@
Figure 6: Australian Bushfire Heatwave: (a) Testing the justification for a GEV. The red curve indicates the empirical CDF of the highest annual 7-day running mean of daily maximum temperature over the area defined in van Oldenborgh et al. (2021) for observations. The purple and blue lines denote GEV (with a constant mean), and normal distribution fits, respectively. The light and dark green shadings correspond to 90th and 95th percentile confidence intervals. (b) Justification for using a GMST-dependent mean. Box plots of the deviance statistic between a GEV without a GMST component and a GEV with a GMST component using the Australia Bushfire heatwave example with 5,000 bootstrap resamples. The edges of the boxes correspond to the first and third quartiles. The deviance statistic was computed for each of the 35 members from the CESM Large Ensemble (blue) and compared with the 95th percentile significance level from a $\chi^{2}$ distribution with one degree of freedom (red line). A box plot for the deviance statistic for observations is shown in green.
@@
The regression analysis of Australian extreme temperatures and GMST, (Fig. 7a) as well as the range of $\alpha$ values estimated from observations (Fig. 7b; red) both show a GMST dependence near 2 C change in extreme temperatures per degree increase in GMST. As mentioned above, a sensible null hypothesis for heatwaves in a warming climate is that the extreme values shift with the mean warming (Tziperman, 2022). The GMST dependence calculated here suggests that Australian heatwaves increase at nearly twice the rate of the GMST. Such a strong response requires a physical mechanism. Some candidates are (1) the fact that land heats up more quickly than the ocean, so GMST always tends to lag behind land surface temperatures; (2) soil-drying can lead to further increases in warming over land as less water is available for evaporative cooling, etc. It is not clear that these can lead to a doubling of the GMST effect but further exploration of this amplification is beyond the scope. The alternative is that this is a result of internal variability — possibly, ENSO — affecting both Australian heat extremes and the GMST, consistent with Fig. 1a. Again, this suggests that using this value of $\alpha$ for attribution may lead to biased results. Finally, Fig. 7b (blue) shows a large range of GMST-dependencies for different ensemble members in the CESM Large Ensemble, again highlighting that a large component of this dependence most likely represents internal variability rather than anthropogenic change. The distribution of all GEV parameters for the observations is given in Supplementary Fig. SI-4. How the GEV parameters scale with the number of CESM Large Ensemble members used is indicated in Supplementary Fig. SI-5.
@@
Figure 7: Australian Bushfire Heatwave: distribution of $\alpha$ values. (a) Scatter plot of the highest annual 7-day running mean of daily maximum temperatures against GMST (following van Oldenborgh et al., 2021). The red line indicates a linear regression fit to the data, and the blue line indicates a linear fit using the location parameter and $\alpha$ derived from the GEV maximum likelihood estimate. (b; red) PDF of the GEV-fit $\alpha$ distributions from observations, using bootstrap resampling. The $\alpha$ value from the observations without resampling is shown by the red line (at 1.7) (b; blue) Effect of the internal variability and number of data from a model large ensemble on the uncertainty in $\alpha$ for the Australian Bushfire heatwave example. Shown is a histogram of the GEV-fit $\alpha$ PDFs derived from the $\alpha$ value from each ensemble member with no resampling. The average $\alpha$ value across the Large Ensemble is indicated with the blue line.